Grade 11Mechanics

Two forces are such that sum of thier magnitude is 18N, the resultant is √228 when they are at 120. Then the magnitude of the forces are
  1. 12N,6N
  2. 13N,5N
  3. 10N,9N
  4. 16N,2N

Yash
8 Years agoGrade 11

1 Answer

Khimraj
ApprovedApproved Tutor Answer8 Years ago
Let magnitude of vectors are A and B
then
A + B = 18        – (1)
and
\sqrt{228} = \sqrt{A^{2} + B^{2} + 2ABcos120} = \sqrt{A^{2} + B^{2} - AB}
SO 
A2 + B2 – AB = 228               – (2)
BY solving equation (1) and (2)
A = 16 and B = 2
Hope it clears. If you like answer then please approve it.